# Dynamical Moiré Systems in Twisted Bilayer Optical Lattices

### v3

@@ I System and model
$w(x,y)=\sqrt{2\kappa}e^{-\pi\kappa(x^{2}+y^{2})},$ (S14)
@@
$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)},$ (S15)
@@
$\begin{cases}U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa\\ U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa e^{-\pi\kappa d^{2}}\\ U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=&\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)}\end{cases}.$ (S16)
@@
$+\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}.$ (S17)
@@
We investigate the physics of this system by calculating its ground state via the bosonic Gutzwiller variational approach (Krauth et al., 1992; Jaksch et al., 1998; Lanatà et al., 2012). Within the framework of mean-field theory, we use the bosonic Gutzwiller variational approach to investigate the ground state properties of the system, with the variational ground state assumed to be the site-factorized form
$|\mathrm{GW}\rangle=|\phi_{1_{\sigma}}\rangle_{1_{\sigma}}\otimes\ldots\otimes|\phi_{N_{\mathrm{lat},\sigma}}\rangle_{N_{\mathrm{lat},\sigma}}.$ (S18)
@@
$\left.-J\sum_{\mathbf{j}_{\sigma}}\left(\frac{\sum_{n=0}^{\infty}c_{n}^{*(\mathbf{i}_{\sigma})}c_{n+1}^{(\mathbf{i}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}}\right)\left(\frac{\sum_{n=0}^{\infty}c_{n}^{(\mathbf{j}_{\sigma})}c_{n+1}^{*(\mathbf{j}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\sigma})}\right|^{2}}\right)\right]$ (S19)
@@ II Phase diagrams for integer filling
Figure S1: Phase diagrams and typical configurations of the MI-MI case and SF-MI case. (a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ at the filling with $\rho_{\uparrow}=1$ and $\rho_{\downarrow}=1$ (twist angle is fixed at $\theta=36.87^{\circ}$). (b) Similar to (a), but at the filling with $\rho_{\uparrow}=1/5$ and $\rho_{\downarrow}=1$. (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
@@
In the main text, we have focused on the SF-SF case, where both components have non-integer fillings, rather than on the MI-MI case (both components with integer fillings) or the SF-MI case (one component with a non-integer filling and the other with an integer filling). This choice is due to the large excitation gap caused by the strong interactions in the integer-filling component, which hinders atom tunneling and the formation of rich phases, unlike in the SF-SF case. Here, we present the phase diagrams and typical configurations for the MI-MI and SF-MI cases (see Fig. S1). Similar to the conventional Bose-Hubbard model, when the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is small, the components with integer filling favor the homogeneous superfluid phase and Mott-insulator phase [see Figs. S1 (c1) and (c2)], while the component with non-integer filling favor the homogeneous superfluid phase [see Figs. S1 (d1) and (d2)]. In the MI-MI case, even when $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is large, the density distributions of both components remain almost homogeneous due to the excitation gap [see Figs. S1 (c3) and (c4)]. Conversely, in the SF-MI case, a large $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ drives the component with non-integer filling into an induced Mott-insulator phase [see Fig. S1 (d4)]. Moreover, in the regime of relatively large $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ (where interlayer interactions dominate over intralayer interactions), the intense competition between the integer-filling components due to the interlayer interaction tends to make the system more prone to forming a superfluid phase. However, in regions dominated by interlayer interactions, stronger interlayer interactions make it easier to form a Mott-insulator phase [see Fig. S1 (a)]. In the SF-MI scenario, the high-filling component can create an effective potential that localizes the low-filling component, thereby facilitating the formation of the Mott-insulator phase [see Fig. S1 (b)].

### v4

@@ Appendix A Model Hamiltonian and Gutzwiller method
$w(x,y)=\sqrt{2\kappa}e^{-\pi\kappa(x^{2}+y^{2})},$ (13)
@@
$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)},$ (14)
@@
$\begin{cases}U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa\\ U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa e^{-\pi\kappa d^{2}}\\ U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=&\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)}\end{cases}.$ (15)
@@
$+\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}.$ (16)
@@
We investigate the physics of this system by calculating its ground state via the bosonic Gutzwiller variational approach Krauth et al. (1992); Jaksch et al. (1998); Lanatà et al. (2012). Within the framework of mean-field theory, we use the bosonic Gutzwiller variational approach to investigate the ground state properties of the system, with the variational ground state assumed to be the site-factorized form
$|\mathrm{GW}\rangle=|\phi_{1_{\sigma}}\rangle_{1_{\sigma}}\otimes\ldots\otimes|\phi_{N_{\mathrm{lat},\sigma}}\rangle_{N_{\mathrm{lat},\sigma}}.$ (17)
@@
$\left.-J\sum_{\mathbf{j}_{\sigma}}\left(\frac{\sum_{n=0}^{\infty}c_{n}^{*(\mathbf{i}_{\sigma})}c_{n+1}^{(\mathbf{i}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}}\right)\left(\frac{\sum_{n=0}^{\infty}c_{n}^{(\mathbf{j}_{\sigma})}c_{n+1}^{*(\mathbf{j}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\sigma})}\right|^{2}}\right)\right]$ (18)
@@
Figure 4: Phase diagrams and typical configurations of the MI-MI case and SF-MI case. (a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ at the filling with $\rho_{\uparrow}=1$ and $\rho_{\downarrow}=1$ (twist angle is fixed at $\theta=36.87^{\circ}$). (b) Similar to (a), but at the filling with $\rho_{\uparrow}=1/5$ and $\rho_{\downarrow}=1$. (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
Appendix B Phase diagrams for integer filling
In the main text, we have focused on the SF-SF case, where both components have non-integer fillings, rather than on the MI-MI case (both components with integer fillings) or the SF-MI case (one component with a non-integer filling and the other with an integer filling). This choice is due to the large excitation gap caused by the strong interactions in the integer-filling component, which hinders atom tunneling and the formation of rich phases, unlike in the SF-SF case.
@@ Appendix B Phase diagrams for integer filling
Here, we present the phase diagrams and typical configurations for the MI-MI and SF-MI cases (see Fig 4). Similar to the conventional Bose-Hubbard model, when the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is small, the components with integer filling favor the homogeneous superfluid phase and Mott-insulator phase [see Figs 4 (c1) and (c2)], while the component with non-integer filling favors the homogeneous superfluid phase [see Figs 4 (d1) and (d2)]. In the MI-MI case, even when $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is large, the density distributions of both components remain almost homogeneous due to the excitation gap [see Figs 4 (c3) and (c4)]. Conversely, in the SF-MI case, a large $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ drives the component with non-integer filling into an induced Mott-insulator phase [see Fig 4 (d4)]. Additionally, the phase boundary in Fig 4 (a) reflects the system’s evolution as the interlayer interaction increases: from being dominated by intralayer interactions, transitioning to a regime where interlayer and intralayer interactions are in balanced competition, and finally to a regime dominated by interlayer interactions. In the SF-MI scenario, the high-filling component can create an effective potential that localizes the low-filling component, thereby facilitating the formation of the Mott-insulator phase as $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ increases [see Fig 4 (b)].
Figure 5: (a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ with $\rho_{\uparrow}=3/5$ and $\rho_{\downarrow}=3/5$ (the twist angle $\theta$ is fixed at $\theta=7.63^{\circ}$ and the system size is $37\times 37$). (b) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dot in (a). (c) Similar to (a), but the system size is $49\times 49$. (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dot in (c).
Four degenerate ground states
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$E=56.56U$ $E=56.56U$ $E=56.56U$ $E=56.56U$
Metastable States
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$E=60.26U$ $E=58.20U$ $E=57.47U$ $E=56.72U$
Table 1: Degenerate ground state and metastable state configurations for the system parameter that corresponds to Fig. 3(d2) in the manuscript.
Appendix C Finite-size effects
In this section, we present a detailed analysis of the finite-size effects in our numerical simulations. Given the spatial dependence of the interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ and the potential impact of the system size on the observed physical properties, we have carried out additional simulations for larger lattice sizes to ensure the reliability of our results.
To investigate the finite-size effects, we compare the phase diagrams and typical physical quantities between two system sizes: $37\times 37$ and $49\times 49$. Specifically, we focus on the smallest twist angle $\theta=7.63^{\circ}$ studied in our manuscript. Figs 5 (a) and (c) shows a direct comparison of the phase diagrams for the $37\times 37$ and $49\times 49$ systems, and both phase diagrams display qualitatively consistent results. In addition to the phase diagrams, we also compare the spatial configurations of relevant physical observables, such as the superfluid order parameter $\phi_{\mathbf{i}_{\sigma}}\equiv\langle\hat{b}_{\mathbf{i}_{\sigma}}\rangle$ and density distribution $\langle\hat{n}_{\mathbf{i}_{\sigma}}\rangle$. Figs 5 (b) and (d) display representative configurations for both system sizes. The spatial profiles of the superfluid order parameter and density exhibit negligible differences between the two system sizes, further confirming that the 37×37 system adequately captures the essential physics of the system.
Appendix D Spontaneously breaking of $C_{4}$ symmetry and metastable states
In this section, we investigate the spontaneous breaking of $C_{4}$ rotational symmetry and the metastable states observed in our simulations. These phenomena arise due to the intrinsic disorder potential, leading to the four degenerate ground states and a rich set of metastable states. To demonstrate the degeneracy of these ground states, we performed simulations with different initial conditions for the system parameters corresponding to Fig. 3(d2) of the main text. As shown in the first row of the plots in Table 1, we identified four distinct ground state configurations possessing identical energy. This confirms the spontaneous breaking of the $C_{4}$ symmetry in the system.
In addition to the four degenerate ground states, our simulations reveal the presence of numerous metastable configurations within the Bose glass phase. These metastable states have higher energies than the ground states. The second row of the plots in Table 1 provides a representative set of metastable configurations for the same parameter, illustrating the diversity of particle distributions and their associated energies. This observation aligns with the expected behavior of a Bose glass, which is typically characterized by the coexistence of multiple metastable configurations.
Figure 6: Identification of the Bose glass phase. (a) Typical configurations of the Bose glass phase [induced Mott insulator phase with non-percolated superfluid islands as well as the Fig. 2(c4) in the main text]. (b) Typical configurations of the induced Mott insulator phase with percolated superfluid region [Fig. 2(c3) in the main text]. (c) Typical configurations of the superfluid phase [Fig. 2(c2) in the main text]. The percolated superfluid regions across the length scale of the system size (the red square) are marked by the small red circles.
Appendix E Bose glass identification
To identify the Bose glass phase in our numerical simulations, we implemented a systematic approach based on the percolation properties of the superfluid regions and the spatial characteristics of the density distribution. This section outlines the methodology used to identify the Bose glass phase.
We use the approach outlined by references Niederle and Rieger (2013); Johnstone et al. (2021b) to analyze percolation of SF clusters on a discrete function $S_{\mathbf{i}_{\sigma}}$. We label the lattice site $\mathbf{i}_{\sigma}$ with integer density distribution $\langle\hat{n}_{\mathbf{i}_{\sigma}}\rangle$ and zero SF order parameter $\phi_{\mathbf{i}_{\sigma}}$ as MI site with $S_{\mathbf{i}_{\sigma}}=0$, and label the site with non-integer filling and nonzero $\phi_{\mathbf{i}_{\sigma}}$ as SF site with $S_{\mathbf{i}_{\sigma}}=1$. The BG phase is identified by examining whether a percolated superfluid region exists on the length scale of the system size. If none of the connected clusters formed by the SF sites ($S_{\mathbf{i}_{\sigma}}=1$) percolates across the system, the phase is identified as the BG phase. This phase is characterized by isolated SF clusters embedded within an MI sea ($S_{\mathbf{i}_{\sigma}}=0$), resulting in a non-vanishing density of SF sites that lack global connectivity Niederle and Rieger (2013). Indeed, in the typical BG phase observed in our numerical simulations, isolated superfluid islands (i.e., non-percolated superfluid regions) are usually surrounded by MI regions. The “induced MI” phase defined in the main text is characterized by regions of the system where sites are filled with an integer number of atoms and have no associated SF order parameter. Consequently, the BG phase can still be considered a form of “induced MI” as long as there are MI regions present.
Following the percolation identification approach on a square lattice described in references Niederle and Rieger (2013); Barman et al. (2013), we apply a similar method in our system with an external cylinder box trap. Specifically, we inscribe the largest possible square within the circular region to assess percolation. To ensure robustness, this inscribed square is systematically rotated from $0{{}^{\circ}}$ to $360{{}^{\circ}}$, allowing us to evaluate percolation for all orientations. Examples of the identification of BG phase are presented Fig. 6.